# Explanation Regularisation through the Lens of Attributions

### v1

@@ Acknowledgements
Appendix A Attention-Rollout
Rollout was introduced in Abnar and Zuidema (2020), and makes it possible to obtain global attributions by recursively aggregating vector-based local attributions, such as attention. We can also use it to show why we expect global attributions to be more impactful than the local counterparts when supervising a model with explanations. By defining attention-rollout recursively, $\mathbf{a}_{l}=\rho(\mathbf{h}^{l},\mathbf{h}^{l-1};\mathbf{a}_{l-1})$, where $\mathbf{a}_{l}$ corresponds to attention-rollout weights at layer $l$, and $\mathbf{h}^{l}$ to the hidden states at layer $l$, we can write its gradient. Namely, for the $l$th layer, we get:
@@ Appendix A Attention-Rollout
By inspecting Equation 5, we can observe how the impact of guiding with a global vector-based technique goes beyond that of a local technique – there we have $\mathbf{a}_{l}=\rho(\mathbf{h}^{l},\mathbf{h}^{l-1})$, meaning that the last recursion term of the gradient that propagates to the layers below would not be part of the computation. This difference seems to indicate that using a global attribution technique, such as attention-rollout, as $E$ in Equation 3 will more strongly limit the model’s ability to condition on features other than the rationale tokens. We show a visual interpretation of this difference in Figure 5.
@@
Appendix B Plausibility Metrics
In order to measure how well the input attributions match the human annotated tokens we employ three metrics introduced in Fomicheva et al. (2021).
@@ Appendix B Plausibility Metrics
As mentioned in Fomicheva et al. (2021), AUC scores might be optimistic when dealing with unbalanced data. That is relevant when working with rationales, where the annotated tokens might correspond to a small portion of the input. Thus, we report the average precision score, which summarizes the average-precision curve as a weighted average of the precisions at the different thresholds,
@@
Appendix C AUC Scores per Layer
As a complement to the results in Sections 6.2-6.4, Figure 6 shows average plausibility AUC scores across layers for baseline, ER with attention as the guided attribution technique (ER+Att and ER-C+Att), and ER with attention-rollout as the guided attribution technique (ER+AttR and ER-C+AttR). We can observe that the increase in plausibility for attention at the top-layer for ER+Att and ER-C+Att is not representative of changes in the plausibility of the model attributions. This observation holds for the attention attributions, and also for the ones obtained with ALTI and DecompX. On the other hand, attention-rollout (in the constrained setting) leads to changes in attributions for all layers, adding further evidence of its ability to impact the ER model to rely on plausible tokens.
Figure 6: Average AUC plausibility scores per-layer ($\uparrow$) computed with Attention, ALTI, and DecompX.
Appendix D Predicting OOD Performance
@@
To complement the results discussed in Section 6.6, we obtain results on a wider range of OOD datasets and plausibility metrics criteria, and include also joint ER+Att and ER+AttR results. Results are shown in Figure 7. Similarly to Section 6.6, no in-domain criteria exhibits correlation with OOD classification performance.
(a) SST-Dev Classification as Predictor
(b) SST-Dev Plausibility as Predictor - AUC with Attention
(c) SST-Dev Plausibility as Predictor - AUC with Attention-Rollout
(d) SST-Dev Plausibility as Predictor - AUC with DecompX-Classifier
Appendix E Extra OOD Annotations
@@
In order to further support our claims with respect to increased robustness to OOD conditions, we annotate 50 examples of the Yelp dataset. We first filter out examples that will result in more than 150 subtokens when using the tokenizer from the pre-trained encoder model we use during training (so that we can compute attributions with all available metrics, in particular ALTI (Ferrando et al., 2022) and DecompX (Modarressi et al., 2023), which we find to result in prohibitive runtime or OOM errors with long sequences). Then, we sample 55 reviews and label pairs at random from those, tokenize these examples with the English tokenizer from spaCy (Honnibal et al., 2020), and manually annotate the first five, selecting all sequences of text that offer evidence for the gold label. Those are use as few-shot examples for a prompt to Meta-Llama-3-8B-Instruct [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct (AI@Meta, 2024) that outputs automatic annotations for the remaining 50 examples. Finally, we post-edit the automatic annotations whenever necessary.
@@
Table 5: F1-Macro scores ($\uparrow$) for the Yelp-50 split used to compute OOD attributions.
Appendix F Hyperparameters and Training
Hyperparameters.
We choose the combination of learning rate and maximum number of training epochs (which is relevant since we use a linear scheduler with 10% of training steps as warm-up) with the lowest cross-entropy loss over three runs with different seeds. For the explanation regularised approach we explore the same hyperparameters, plus the $\lambda$ weight. We select based on cross-entropy loss so that different values of $\lambda$ can be directly compared.
For the constrained approach, we first train a model (3 seeds) using $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$ and use the (rounded) average of the obtained minimum explanation train and validation losses as our bounds $b_{\mathrm{train}}$ and $b_{\mathrm{val}}$. Then, we choose the combination of learning rate and constrained optimizer learning rate that minimizes the average cross entropy loss.
Training.
For constrained optimization, we choose the model checkpoint with the lowest validation cross entropy loss, provided that the validation explanation loss is $\mathcal{L}_{\mathrm{expl}_{\mathrm{val}}}<1.1\times b_{\mathrm{val}}$. This ensures that we compare model checkpoints where the guided attribution technique is learning to predict the annotated rationales, according to the defined bound. Selecting a checkpoint based on the same criteria would have not been possible for the joint approach, as none of the runs converges to an explanation loss value that meets the defined bound. This further highlights the benefits of constrained optimization as analysis tool for our use case.
Unless mentioned otherwise, we report results over 15 seeds. All our experiments are developed using a single Nvidia A100 40GB GPU, and implemented with PyTorch (Paszke et al., 2019) and Pytorch Lightning (Falcon and The PyTorch Lightning team, 2019).
F.1 Hyperparameters
For the baseline we explore learning rate $\in\{2,3,5\times 10^{-5}\}$ and maximum number of epochs $\in\{15,25\}$. For the joint approach we explore the same space as the baseline model and $\lambda\in\{0.6,1.0,1.4\}$, following a set of choices aligned with previous work on ER for OOD robustness. For the constrained approach we explore learning rate $\in\{2,3,5\times 10^{-5}\}$ and constrained learning rate $\in\{1\times 10^{-1},5\times 10^{-2}\}$.
Final choices can be seen in Table 6. The trained models for Section 6.5 all use the same hyperparameters as the baseline model.
Shared Value
Batch Size 32
Optimizer AdamW
Weight Decay 0
$\beta$ (0.9, 0.98)
Scheduler Linear
Scheduler Warmup Steps 10%
@@ F.1 Hyperparameters
$b_{\mathrm{train}}$ – Initial 0.035
$b_{\mathrm{train}}$ – Minimum 0.023
@@
$b_{\mathrm{train}}$ – Initial 0.03
$b_{\mathrm{train}}$ – Minimum 0.022
@@
Table 6: Hyperparameters. (Loshchilov and Hutter, 2018)
F.2 Data
For in-domain and training data we use SST-2 (Socher et al., 2013), following the heuristic algorithm proposed in Carton et al. (2020) to obtain instance-level rationales [^9] https://github.com/BoulderDS/evaluating-human-rationales/blob/master/scripts/download_and_process_sst.py. For out-of-domain we use: Amazon-Reviews [^10] https://huggingface.co/datasets/McAuley-Lab/Amazon-Reviews-2023 (Hou et al., 2024), IMBD [^11] https://huggingface.co/datasets/stanfordnlp/imdb (Maas et al., 2011), Movies [^12] https://huggingface.co/datasets/eraser-benchmark/movie_rationales (Zaidan and Eisner, 2008; DeYoung et al., 2020), and Yelp [^13] https://huggingface.co/datasets/fancyzhx/yelp_polarity (Zhang et al., 2015). For some of the OOD evaluation data, the full test splits contain several thousand examples. In those cases we sample 5,000 examples. More details on the final data are described in Table 7. All data is in English.
Train Dev Test % Positive Avg. / Max Length Source Rationales
IMDB (Maas et al., 2011) - - 5,000 50.7 % 288 / 1024 Movies
$\times$
$\times$
Yelp (Zhang et al., 2015) - - 5,000 50.6 % 168 / 1024 Businesses
$\times$
Yelp-50∗∗ - - 50 50.0 % 84 / 151 Businesses
Table 7: Sentiment analysis datasets, including percentage of examples labeled as positive, average and maximum length of inputs (after truncating), source of the data, and whether rationales are provided. – This includes both dev and test examples. ∗∗ – This corresponds to the subset of annotated data described in Section E.
F.3 Guided Attribution Technique
@@
Following Joshi et al. (2022), we scale the attributions $E(C_{\theta},x)$ by 100 when computing the explanation loss (Equation 3) and re-normalize them with softmax. We take it as part of the approach, and do so for both approaches that use attention as the guided attribution technique, ER+Att and ER-C+Att.
F.4 Evaluation Attribution Techniques
We keep all default choices part of the original work for DecompX, DecompX-C, and ALTI. For post-hoc attributions techniques that output ‘signed’ attribution scores in Section 5.1, i.e., IxG and DX-C, we take the absolute value of the attribution score.
F.5 Classifier
@@
Following Joshi et al. (2022), classification is performed with a pre-trained Transformer encoder model, google/bigbird-roberta-base [^14] https://huggingface.co/google/bigbird-roberta-base (Zaheer et al., 2020). This is followed by a linear layer that uses as input the top-layer representation of the [CLS] token, with tanh as the non-linearity, and no classifier dropout.
@@
Att
AttR
IxG
ALTI
DX-C
DX
Att
AttR
IxG
ALTI
DX-C
DX
Att
AttR
IxG
ALTI
DX-C
DX
@@
Table 8: Attributions to Annotated Tokens. AUC, Average Precision and Recall@k Scores ($\uparrow$) for SST-Dev (ID), Movies (OOD), and Yelp-50 (OOD) datasets, for the Baseline (Baseline), ER model guided with Attention (ER + Att), ER model guided with Attention-Rollout (ER + AttR), and the corresponding constrained counterparts (ER-C + X). Attributions are computed using Attention (Att), Attention + Rollout (AttR), InputxGradient (IxG), ALTI Aggregated (ALTI), and DecompX (with) Classifier (DX and DX-C) as the attribution techniques. For Movies, we report only the techniques that can be applied to longer sequences.

### v3

@@ Acknowledgements
Appendix A Attribution Techniques
We use a total of six attribution techniques in this work. They serve two main roles, acting as: (i) guided attribution techniques, when used as part of ER training to compute the explanation loss, $\mathcal{L}_{\mathrm{expl}}$; and (ii) non-guided attribution techniques, when used to evaluate the impact of ER on input attributions. The attribution techniques we use are either gradient-based or vector-based. In the case of the latter, they differ on the components of the Transformer encoder architecture (Vaswani et al., 2017) that are used, and also on multiple design choices. Finally, all techniques provide global attributions, with the exception of attention, which is a local attribution technique.
Attention.
Attention attributions correspond to the normalised top-layer attention weights. These are directly obtained from the Transformer model architecture. We use the values from the [CLS] token, averaged across heads.
Attention-Rollout.
Attention-Rollout (Abnar and Zuidema, 2020) attributions are obtained by recursively aggregating attention scores. As in the original implementation, we incorporate the residual connection by adding the identity matrix ($\mathbf{I}$) to the original attention matrix ($\mathbf{A}^{l}$), followed by a normalization step, resulting in $\mathbf{A}^{l}_{R}=0.5\mathbf{A}^{l}+0.5\mathbf{I}$. We use the values from the [CLS] token, averaged across heads.
InputXGradient.
InputXGradient (Shrikumar et al., 2017) attributions are obtained by multiplying the input value by the corresponding gradient with respect to the output target label. We use the Captum package implementation (Kokhlikyan et al., 2020).
ALTI.
ALTI (Ferrando et al., 2022) is a vector-based attribution technique based on the attention block decomposition introduced in Kobayashi et al. (2021). This decomposition factors in not only the attention weights, but also the value vectors, the output projection, as well as the first residual connection and layer normalization. The main difference lies in how the value of the contribution of a token at the layer level is computed — ALTI takes into account the output vector of the attention block and uses the L1 norm instead of the L2 norm. Furthermore, ALTI then aggregates the local attributions over the full model using a rollout-like approach, resulting in global attributions. We adapt the available code. [^5] https://github.com/mt-upc/transformer-contributions
DecompX.
DecompX (Modarressi et al., 2023), similarly to ALTI, is a vector-based attribution technique based on the attention block decomposition by Kobayashi et al. (2021). The main differences are twofold: first, DecompX incorporates all the Transformer encoder components, including the feed-forward networks. Second, instead of using a recursive approach like rollout to aggregate the local attributions, DecompX propagates ‘decomposed token representations’ through the model. We also report DecompX-Classifier, which includes the classification head, and whose output is signed. We adapt the available code. [^6] https://github.com/mohsenfayyaz/DecompX
Appendix B Impact of Local vs Global Guided Attributions in ER
We can use attention-rollout to show why we expect global attributions to be more impactful than the local counterparts when supervising a model with explanations. Rollout was introduced in Abnar and Zuidema (2020), and makes it possible to obtain global attributions by recursively aggregating vector-based local attributions, such as attention. By defining attention-rollout recursively, $\mathbf{a}_{l}=\rho(\mathbf{h}^{l},\mathbf{h}^{l-1};\mathbf{a}_{l-1})$, where $\mathbf{a}_{l}$ corresponds to attention-rollout weights at layer $l$, and $\mathbf{h}^{l}$ to the hidden states at layer $l$, we can write its gradient. Namely, for the $l$th layer, we get:
@@ Appendix B Impact of Local vs Global Guided Attributions in ER
By inspecting Equation 3, we can observe how the impact of guiding with a global vector-based technique goes beyond that of a local technique – there we have $\mathbf{a}_{l}=\rho(\mathbf{h}^{l},\mathbf{h}^{l-1})$, meaning that the last recursion term of the gradient that propagates to the layers below would not be part of the computation. This difference seems to indicate that using a global attribution technique, such as attention-rollout, as $E$ in the explanation loss will more strongly limit the model’s ability to condition on features other than the rationale tokens. We show a visual interpretation of this difference in Figure 7.
@@
Appendix C Attributions Techniques Faithfulness
To further validate our choice of attribution techniques, we assess the faithfulness of the used input attributions by computing normalised sufficiency and comprehensiveness (Carton et al., 2020). These are computed as follows:
$\mathrm{NullDiff}(x,\hat{y})=\mathrm{max}(0,p(\hat{y}|x)-p(\hat{y}|x,0))$ (4)
$\mathrm{NormSuff}(x,\hat{y},\alpha)=\frac{\mathrm{Suff}(x,\hat{y},\alpha)-\mathrm{Suff}(x,\hat{y},0)}{1-\mathrm{Suff}(x,\hat{y},\alpha)}$ (5)
$\mathrm{NormComp}(x,\hat{y},\alpha)=\frac{\mathrm{Comp}(x,\hat{y},\alpha)}{\mathrm{Comp}(x,\hat{y},1)},$ (6)
with $\hat{y}=\mathrm{arg~{}max}~{}p(y|x)$. Both normalised sufficiency and comprehensiveness, computed with the top 1%, 20%, 40%, 60%, 80%, and 100% most attributed tokens, can be seen in Figure 8. As expected, both metrics improve the more top-attributed tokens, based on a given attribution technique, are used as input to the classifier. Furthermore, we find DecompX, DecompX-Classifier, and ALTI to perform the best. This corroborates the results reported on their respective works.
Figure 8: Normalised sufficiency and comprehensiveness ($\uparrow$) scores for the top-xtokens according to a given attribution technique, for SST-Dev, using the Baseline approach.
Appendix D Plausibility Metrics
We employ three metrics introduced in Fomicheva et al. (2021, 2022) to measure how well the input attributions match the human annotated tokens.
@@ Appendix D Plausibility Metrics
As mentioned in Fomicheva et al. (2021), AUC scores might be optimistic when dealing with unbalanced data. That is relevant when working with rationales, where the annotated tokens might correspond to a small portion of the input. Thus, we report the average precision score, which summarizes the average-precision curve as a weighted average of the precision scores at the different thresholds,
@@
Appendix E Input Attributions Correlations
As discussed in Section 5, one of the tools we use to study the impact of ER on the input model attributions is to measure changes in the correlation of attribution scores for a fixed technique across approaches, e.g., how does attention correlate baseline vs baseline and baseline vs ER+Att. Here, for each attribution technique, we iterate over all examples, sample two model versions, select the corresponding attributions for the Baseline and approach we want to assess, and then compute the respective Kendall rank correlation coefficient [^7] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.kendalltau.html. We show results for SST-Dev, Yelp-50, and Movies using all ER approaches and attribution techniques in Figure 9. The conclusions are as discussed in Section 5: ER techniques based on attention (ER+Att and ER-C+Att) impact only the attention attributions, while global constrained ER approaches (ER-C+AttR and ER-C+IxG) are able to impact all reported attribution techniques.
Alternatively, it is also possible to compare correlation coefficients for a pair of attribution techniques for a given choice of training objective (e.g., the correlation between attention and attention-rollout for baseline or ER+Att) and how those correlations change as we vary the training objective. We compute average Kendall rank correlation values and report them in Figure 10. The obtained results agree with our previous findings: (i) attention as a guided attribution technique (ER+Att and ER-C+Att) impacts mostly only attention attributions; (ii) global guided attribution techniques have little impact when used in joint ER (ER+AttR and ER+IxG); (iii) only the constrained ER approaches with global guided attributions techniques are able to clearly impact most attributions; and (iv) for OOD most patterns are the same, despite less noticeable. Note how this correlation analysis is different from the previous. First, here we inspect averages (the number of data points makes it unfeasible to report histograms of correlation coefficients), which has the potential to be misleading. Second, we correlate across attribution techniques instead of correlating the same attribution across approaches, which leads to observations that are more complex and difficult to interpret.
Appendix F Predicting OOD Performance
@@
To complement the results discussed in Section 5.3, we present findings on a broader range of OOD datasets and plausibility metrics criteria, also including joint ER+Att and ER+AttR results. Results are shown in Figure 14. Similarly to Section 5.3, no in-domain criteria exhibits correlation with OOD classification performance.
Appendix G Data
For in-domain and training data we use SST-2 (Socher et al., 2013), following the heuristic algorithm proposed in Carton et al. (2020) to obtain instance-level rationales [^8] https://github.com/BoulderDS/evaluating-human-rationales/blob/master/scripts/download_and_process_sst.py. For out-of-domain we use: Amazon-Reviews [^9] https://huggingface.co/datasets/McAuley-Lab/Amazon-Reviews-2023 (Hou et al., 2024), IMBD [^10] https://huggingface.co/datasets/stanfordnlp/imdb (Maas et al., 2011), Movies [^11] https://huggingface.co/datasets/eraser-benchmark/movie_rationales (Zaidan and Eisner, 2008; DeYoung et al., 2020), and Yelp [^12] https://huggingface.co/datasets/fancyzhx/yelp_polarity (Zhang et al., 2015). From these OOD datasets only Movies includes human annotated rationales. Some OOD evaluation datasets have several thousand examples. In those cases we sample 5,000 examples. More details on the final data are described in Table 6. All data is in English.
Train Dev Test % Positive Avg. / Max Length Source Rationales
IMDB (Maas et al., 2011) - - 5,000 50.7 % 288 / 1024 Movies
$\times$
$\times$
Yelp (Zhang et al., 2015) - - 5,000 50.6 % 168 / 1024 Businesses
$\times$
Yelp-50∗∗ - - 50 50.0 % 84 / 151 Businesses
Table 6: Sentiment analysis datasets, including percentage of examples labeled as positive, average and maximum length of inputs (after truncating), source of the data, and whether rationales are provided. – This includes both dev and test examples. ∗∗ – This corresponds to the subset of annotated data described in Section G.1.
G.1 Extra OOD Annotated Data
@@
We conduct our OOD analysis on the Movies dataset, an available sentiment analysis dataset with human rationale annotations. To further support our claims with respect to increased robustness to OOD conditions, we annotated a split of 50 examples of the Yelp dataset. We first filter out examples that will result in more than 150 subtokens when using the tokenizer from the pre-trained encoder model we use during training (so that we can compute attributions with all available metrics, in particular ALTI (Ferrando et al., 2022) and DecompX (Modarressi et al., 2023), which we find to result in prohibitive runtime or OOM errors with long sequences). Then, we sample 55 examples at random from those, tokenize them using the English tokenizer from spaCy (Honnibal et al., 2020), and manually annotate the first five, selecting all sequences of text that offer evidence for the gold label. These are used as few-shot examples for a prompt to Meta-Llama-3-8B-Instruct [^13] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct (AI@Meta, 2024) that outputs automatic annotations for the remaining 50 examples. Finally, we post-edit the automatic annotations whenever necessary.
@@
ER-C+IxG 94.51 $\pm$ 3.17
Table 7: F1-Macro scores ($\uparrow$) for Yelp-50.
Appendix H Hyperparameters and Training
Optimizer and Scheduler.
We use AdamW (Loshchilov and Hutter, 2018) as our optimizer, with $\beta$ values set to (0.9, 0.98) and the weight decay coefficient set to 0. For the learning rate we use a linear scheduler, with 10% of training steps as warm-up. For constrained optimisation parameters we use RMSprop (Tieleman and Hinton, 2012).
Hyperparameter Selection.
We choose the combination of learning rate and maximum number of training epochs (since we use a linear scheduler with 10% of training steps as warm-up) with the lowest cross-entropy loss over three runs. For the explanation regularised approach we explore the same hyperparameters, plus the $\lambda$ weight. We select based on cross-entropy loss so that different values of $\lambda$ can be directly compared.
For the constrained approach, we first train a model (3 seeds) using $\mathcal{L}=\mathcal{L}_{\mathrm{expl}}$ and use the average minimum explanation train and validation explanation losses to guide our choice of bounds $b_{\mathrm{train}}$ and $b_{\mathrm{val}}$. For $b_{\mathrm{val}}$ we use the average of the minimum validation loss. For $b_{\mathrm{train}}$ we use a value close to 1.5 times the average of the minimum training loss. Then, we choose the combination of learning rate and constrained optimizer learning rate that minimizes the average cross entropy loss.
For the baseline we explore learning rate $\in\{2,3,5\times 10^{-5}\}$ and maximum number of epochs $\in\{15,25\}$. For the joint approach we explore the same space as the baseline model and $\lambda\in\{0.6,1.0,1.4\}$, following a set of choices aligned with previous work on ER for OOD robustness. For the constrained approach we explore learning rate $\in\{2,3,5\times 10^{-5}\}$ and constrained learning rate $\in\{1\times 10^{-1},5\times 10^{-2}\}$. All experiments use a batch size of 32. Final choices can be seen in Table 8. The experiments with multiple lambdas uses the same choices as the baseline model.
@@ Appendix H Hyperparameters and Training
Joint IxG Value
Learning Rate $2\times 10^{-5}$
Train Epochs 25
$\mathbf{\lambda}$ 1.0
@@
$b_{\mathrm{train}}$ 0.035
@@
$b_{\mathrm{train}}$ 0.030
@@
Constrained IxG Value
Learning Rate $2\times 10^{-5}$
Train Epochs 25
Constrained Learning Rate $1\times 10^{-1}$
$b_{\mathrm{train}}$ 0.35
$b_{\mathrm{val}}$ 0.35
Table 8: Hyperparameters choices for all approaches.
Training.
For constrained optimization, we choose the model checkpoint with the lowest validation cross-entropy loss, provided that the validation explanation loss is $\mathcal{L}_{\mathrm{expl}_{\mathrm{val}}}<1.1\times b_{\mathrm{val}}$. This ensures that we compare model checkpoints where the guided attribution technique is learning to predict the annotated rationales, according to the defined bound. Selecting a checkpoint based on the same criteria would have not been possible for the joint approach, as none of the runs converges to an explanation loss value that meets the defined validation bound.
Unless mentioned otherwise, we report results over 15 seeds. [^14] We found ER-C+IxG unstable to train, with only around one in three seeds minimising the CE loss while also minimising the explanation loss. Thus, it required training more models. In practice, we still report results over 15 seeds. All our experiments are developed using a single Nvidia A100 40GB GPU, and implemented with PyTorch (Paszke et al., 2019) and PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019).
Attribution Techniques.
@@
Following Joshi et al. (2022), we scale the attributions $E(C_{\theta},x)$ by 100 when computing the explanation loss and re-normalize them with softmax. We take it as part of the approaches that use attention as the guided attribution technique, ER+Att and ER-C+Att. For InputXGradient we keep all default choices of the Captum package. The output attributions are aggregated into token-level attributions via sum, and normalised with the L2 norm. For DecompX/DecompX-C [^15] https://github.com/mohsenfayyaz/DecompX, and ALTI [^16] https://github.com/mt-upc/transformer-contributions we keep all default choices part of the original work. For attributions techniques that output ‘signed’ attribution scores, i.e., IxG and DX-C, we take the absolute value of the attribution score.
Classifier.
@@
Following Joshi et al. (2022), we use a pre-trained Transformer encoder model, google/bigbird-roberta-base [^17] https://huggingface.co/google/bigbird-roberta-base (Zaheer et al., 2020), followed by a linear layer that uses as input the top-layer representation of the [CLS] token, with tanh as the non-linearity, and no classifier dropout.
(a) SST-Dev (ID)
(b) Movies (OOD)
(c) Yelp-50 (OOD)
Figure 9: Average Kendall Rank correlation between attribution techniques of the different approaches. The BS (Baseline) and ER (corresponding ER approach) text corresponds to the average correlation values.
(a) SST Dev (ID)
(b) Yelp-50 (OOD)
(c) Movies (OOD)
Figure 10: Average Kendall Rank correlation between attribution techniques for the different approaches.
Figure 11: F1-Macro scores ($\uparrow$). ER-C + Att uses attention as the guided attribution technique, ER-C + AttR uses attention-rollout, and ER-C + IxG uses InputXGradient. Results correspond to 15 seeds.
Figure 12: Cross-Entropy vs Explanation Loss for SST-Dev, using ER+IxG. The vertical line shows the validation explanation-loss bound. Each point corresponds to the average of 5 runs, and the error bars to the standard deviation for each loss. $\lambda$ ranges from 0 (No-ER) to 100.
(a) SST-Dev (ID)
(b) Yelp-50 (OOD)
(c) Movies (OOD)
Figure 13: Average AUC plausibility scores per-layer ($\uparrow$). We report only techniques that output per-layer attributions.
(a) SST-Dev Classification as Predictor
(b) SST-Dev Plausibility as Predictor - AUC with Attention
(c) SST-Dev Plausibility as Predictor - AUC with Attention-Rollout
(d) SST-Dev Plausibility as Predictor - AUC with DecompX-Classifier
@@
Att AttR IxG ALTI DX-C DX Att AttR IxG ALTI DX-C DX Att AttR IxG ALTI DX-C DX
@@
$\mathcal{L}_{\text{expl}}$ (IxG) 55.2 $\pm$ 12 58.3 $\pm$ 8.3 68.6 $\pm$ 0.7 72.2 $\pm$ 1.7 58.8 $\pm$ 1.6 60.7 $\pm$ 6.0 37.7 $\pm$ 10 33.4 $\pm$ 7.5 49.5 $\pm$ 0.2 53.6 $\pm$ 2.2 38.2 $\pm$ 2.0 41.1 $\pm$ 8.1 31.6 $\pm$ 11 32.2 $\pm$ 10 44.9 $\pm$ 1.0 47.1 $\pm$ 2.1 34.9 $\pm$ 2.5 38.3 $\pm$ 7.4
Table 9: Attributions to Annotated Tokens. AUC, Average Precision and Recall@k Scores ($\uparrow$) for SST-Dev (ID), Movies (OOD), and Yelp-50 (OOD) datasets, for the Baseline (Baseline), ER model guided with Attention (ER + Att), ER model guided with Attention-Rollout (ER + AttR), ER model guided with InputXGradient (ER + IxG) and the corresponding constrained counterparts (ER-C + X). Attributions are computed using Attention (Att), Attention + Rollout (AttR), InputXGradient (IxG), ALTI Aggregated (ALTI), and DecompX (with) Classifier (DX and DX-C) as the attribution techniques. For Movies, we report only the techniques that can be applied to longer sequences.